### v2

@@ 7.9 Measuring the entanglement
and the entanglement spectrum can be obtained from the SVD decomposition of the $\mathbf{M}$ matrix. Like the Hamiltonian, the $\mathbf{M}$ matrix is block diagonal, and each block carries different quantum numbers of the Hilbert spaces of $A$ and $B$ subsystem [^18] The $\alpha_{o}$ and $M_{IJ}$ in our convention is equivalent to $\mathscr{F}_{m,A}$ and $R_{\mu\nu}^{A}$ in the convention of Ref. [58], the conversions are $\alpha_{mf}=\sqrt{\mathscr{F}_{m,A}}$ and $M_{IJ}=R_{\mu\nu}^{A}/\sqrt{p}$..
@@
we then specify the sectors to calculate : The number of electrons in subsystem $A$ run from 00 to $N_{m}$ ; the angular momenta in subsystem $A$ can take all permitted values [^19] The sectors with no allowed configurations are automatically ignored ; for subsystem $B$, $N_{e,B}=N_{m}-N_{e,A}$, $L_{z,B}=-L_{z,A}$ ; the $\mathbb{Z}_{2}$ sectors of the two subsystems are the same.
@@ 7.11 FuzzyManifolds — module for other geometries
  • The orbitals are now labelled by a $\mathbb{Z}_{N_{m}}$ number $m=1,\dots,N_{m}$, and $m$ and $m+N_{m}$ are identified. Correspondingly, the conservation of angular momentum $L^{z}$ is now modified into a $\mathbb{Z}_{N_{m}}$ translation symmetry with $\sum_{mf}mN_{mf}\mod N_{m}$ conserved. The GetTorusLz2QNDiag function constructs the corresponding QNDiag.
  • The interacting potentials are still parametrised by the pseudopotentials in a similar sense, but the matrix $C_{m_{1}m_{2}m_{3}m_{4}}^{l}$ is different. The functions GetTorusDenIntTerms and GetTorusPairIntTerms generate the density-density interaction terms and the pair-pair interaction terms from the pseudopotentials.
The detail of the setup is given in Appendix B.
@@ 8.1 DMRG with ITensor
In the example of the Ising model, for convenience, we exchange the Pauli matrices $\sigma^{x}$ and $\sigma^{z}$ so that the two flavours carry $\mathbb{Z}_{2}$-charge 00 and $1$. The sites can be constructed by
@@
To calculate the $\mathbb{Z}_{2}$-even $L^{z}=0$ sector, the initial state can be taken as all the $\mathbb{Z}_{2}$-even sites being filled and all the $\mathbb{Z}_{2}$-odd sites being empty [^20] Note that ITensor takes the string "1" instead of the number 1 as occupied and "0" instead of 0 as filled..
@@
The inner product can be measured by the ITensor function inner. For example, to measure the angular momentum $L^{2}$ of the ground state,
@@
8.2 The EasySweep extension
The extension EasySweep facilitates the management of the DMRG process. It automatically records the intermediate results and recover the run from these intermediate results if a job is stopped on HPC. It also manages the gradual increase of maximal bond dimensions and the determination of convergence by the criteria of energy. This extension contains the following functions :
  • GetMPOSites returns the MPO and sites for given operator terms and a Hilbert space with given quantum numbers. The function first checks if the MPO and sites are already stored in a specified file. If they are already stored, they are read and returned. Otherwise, the sites are generated from the quantum numbers, and the MPO is generated from the terms. The MPO and sites are then written into the file and returned [^22] We recommend the package ITensorMPOConstruction stored at https://github.com/ITensor/ITensorMPOConstruction.jl to construct MPOs..
@@ 8.2 The EasySweep extension
To use this extension for ITensor, one needs to use the packages ITensors, ITensorMPS and HDF5. As the setup, a path needs to be created to store the result HDF5 files, and a method that converts OpSum and sites to MPO needs to be defined.
@@ 9 Practical examples
FuzzifiED can help reproduce almost all the ED and DMRG results in fuzzy sphere works. We offer a series of such examples to help the users get started with the package. The code can be found in the examples directory of the source code repository.
@@
    defect_spectrum.jl calculates the spectrum of magnetic line defect in the 3d Ising model in $l_{z}=0,\mathscr{P}=\pm 1$ and $l_{z}=1$ sectors, calibrated by bulk $T$. This example reproduces Table I in Ref. [6].
@@
    defect_correlator.jl calculates the 1-pt function $\sigma$ and 2-pt function $\sigma\hat{\phi}$ of magnetic line defect in the 3d Ising model. The normalisation of the correlators requires extra bulk data. This example reproduces Figure 4 in Ref. [6].
@@
    defect_changing.jl calculates the spectrum of the defect-creation and changing operators of the magnetic line defect in the 3d Ising model. This example reproduces Table 2 and Figure 5 in Ref. [9].
@@
    defect_overlap.jl calculates the $g$-function of magnetic line defect in the 3d Ising model using the ovelaps between the bulk, defect ground state and the lowest defect-creation state. This example reproduces Figure 6 in Ref. [9].
@@
    cusp_dim.jl calculates the scaling dimension of the cusp of the magnetic line defect in the 3d Ising model as a function of the angle $\theta$. This example reproduces Table 2, upper panel in Ref. [11].
@@
    o3_wf_spectrum.jl calculates the spectrum of $\mathrm{O}(3)$ Wilson-Fisher CFT using the bilayer Heisenberg model. This example reproduces Table I and Figure 2 in Ref. [8].
@@ Acknowledgements.
We extend our deepest gratitude to Yin-Chen He, who provided meticulous testing, unwavering support, and invaluable guidance throughout the development of the package. We also thank Liangdong Hu, Wenhan Guo, Andrew Liam Fitzpatrick, Christian Voinea, Xiang Li, and Wei Zhu for their suggestions and contributions to the package. Z. Z. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ A.1 Construction of Lin table
[…]
@@
[…]
@@ A.2 Compressed sparse column (CSC) sparse matrix
In the CSC format, the matrix is stored in three arrays. We first index elements from $1$ to the number of elements $N_{\textrm{el}}$ in the ascending order from the column to column, and in each column from row to row.
@@
[…]
@@
    The array $rowid$ records the row index of each element and is of length $N_{\textrm{el}}$.
@@
[…]
@@
[…]
@@ A.3 Indexing the boson configurations
[…]
@@
[…]
@@
[…]
Appendix B Landau levels and pseudopotentials on the torus
In this appendix, we describe the construction of Landau levels and pseudopotentials on the torus [98, 99, 100], mentioned in Sections 4.4 and 7.11.
@@ B.1 Lowest Landau level
The torus is nothing more than a $l_{x}\times l_{y}$ rectangle with periodic boundary conditions, i.e., $x\sim x+l_{x}$ and $y\sim y+l_{y}$. We adopt the Landau gauge $A_{x}=0,A_{y}=xB$ and set the magnetic length as the unit length $l_{B}=1$. In this way, the number of orbitals is $N_{m}=l_{x}l_{y}/2\pi$. The lowest Landau level wavefunctions are labelled by the momenta along $y$ direction $m$
@@
$\hat{\psi}_{f}(\mathbf{r})=\hat{\phi}_{m}(\mathbf{r})\hat{c}_{mf},\quad\mathbf{r}=(x,y)$ (91)
@@ B.2 Many-body interaction
$H_{\textrm{int}}=\frac{1}{N_{m}^{2}}\int\mathrm{d}^{2}\mathbf{r}_{1}\mathrm{d}^{2}\mathbf{r}_{2}\,U(\mathbf{r}_{1}-\mathbf{r}_{2})n(\mathbf{r}_{1})n(\mathbf{r}_{2}),\qquad n(\mathbf{r})=\psi^{\dagger}(\mathbf{r})\psi(\mathbf{r})$ (92)
@@
$H_{\textrm{int}}=\sum_{m_{1}m_{2}m_{3}m_{4}}\sum_{\mathbf{q}}\frac{U_{\mathbf{q}}}{2N_{m}}e^{iq_{x}(X_{m_{1}}-X_{m_{3}})}e^{-q^{2}/2}\delta^{\prime}_{t,m_{1}-m_{4}}\delta^{\prime}_{m_{1}+m_{2},m_{3}+m_{4}}c^{\dagger}_{m_{1}}c^{\dagger}_{m_{2}}c_{m_{3}}c_{m_{4}},$ (95)
@@
The interactions can alternatively be organised into pseudopotentials $U_{l}$, describing the energy cost of a pair of electrons with relative angular momentum $l$.
@@
$H_{\textrm{int}}$ $=\sum_{lm_{1}m_{2}m_{3}m_{4}}U_{l}C_{m_{1}m_{2}m_{3}m_{4}}^{l}c^{\dagger}_{m_{1}}c^{\dagger}_{m_{2}}c_{m_{3}}c_{m_{4}}$
@@ 41 Es, sts = dmrg(hmt, sti_m ;
C.4 DMRG with Easy Sweep

### v3

@@ 7.9 Measuring the Entanglement
and the entanglement spectrum can be obtained from the SVD decomposition of the $\mathbf{M}$ matrix. Like the Hamiltonian, the $\mathbf{M}$ matrix is block diagonal, and each block carries different quantum numbers of the Hilbert spaces of $A$ and $B$ subsystem. [^19] The $\alpha_{o}$ and $M_{IJ}$ in our convention is equivalent to $\mathscr{F}_{m,A}$ and $R_{\mu\nu}^{A}$ in the convention of Ref. [74], the conversions are $\alpha_{mf}=\sqrt{\mathscr{F}_{m,A}}$ and $M_{IJ}=R_{\mu\nu}^{A}/\sqrt{p}$.
@@
we then specify the sectors to calculate : The number of electrons in subsystem $A$ run from 0 to $N_{m}$ ; the angular momenta in subsystem $A$ can take all permitted values [^20] The sectors with no allowed configurations are automatically ignored ; for subsystem $B$, $N_{e,B}=N_{m}-N_{e,A}$, $L_{z,B}=-L_{z,A}$ ; the $\mathbb{Z}_{2}$ sectors of the two subsystems are the same.
@@ 7.11 FuzzyManifolds — Module for Other Geometries
  • The orbitals are now labelled by a $\mathbb{Z}_{N_{m}}$ number $m=1,\dots,N_{m}$, and $m$ and $m+N_{m}$ are identified. Correspondingly, the conservation of angular momentum $L^{z}$ is now modified into a $\mathbb{Z}_{N_{m}}$ translation symmetry with $\sum_{mf}mN_{mf}\mod N_{m}$ conserved. The GetTorusLz2QNDiag function implemented the corresponding QNDiag. The translation in the other direction permutes $m$ and is implemented in GetTorusTranslQNOffd.
  • The interacting potentials are still parametrised by the pseudo-potentials in a similar sense, but the matrix $C_{m_{1}m_{2}m_{3}m_{4}}^{l}$ is different. The functions GetTorusDenIntTerms and GetTorusPairIntTerms generate the density-density interaction terms and the pair-pair interaction terms from the pseudo-potentials.
The detail of the set-up is given in Appendix B.
@@ 8.1 DMRG with ITensor
In the example of the Ising model, for convenience, we exchange the Pauli matrices $\sigma^{x}$ and $\sigma^{z}$ so that the two flavours carry $\mathbb{Z}_{2}$-charge 0 and $1$. The sites can be constructed by
@@
To calculate the $\mathbb{Z}_{2}$-even $L^{z}=0$ sector, the initial state can be taken as all the $\mathbb{Z}_{2}$-even sites being filled and all the $\mathbb{Z}_{2}$-odd sites being empty. [^21] Note that ITensor takes the string "1" instead of the number 1 as occupied and "0" instead of 0 as filled.
@@
The inner product can be measured by the ITensor function inner. E. g., to measure the angular momentum $L^{2}$ of the ground state,
@@
8.2 The Easy-Sweep Extension
The extension Easy-Sweep facilitates the management of the DMRG process. It automatically records the intermediate results and recover the run from these intermediate results if a job is stopped on HPC. It also manages the gradual increase of maximal bond dimensions and the determination of convergence by the criteria of energy. This extension contains the following functions :
  • GetMPOSites returns the MPO and sites for given operator terms and a Hilbert space with given quantum numbers. The function first checks if the MPO and sites are already stored in a specified file. If they are already stored, they are read and returned. Otherwise, the sites are generated from the quantum numbers, and the MPO is generated from the terms. The MPO and sites are then written into the file and returned. [^23] We recommend the package ITensorMPOConstruction stored at https://github.com/ITensor/ITensorMPOConstruction.jl to construct MPOs.
@@ 8.2 The Easy-Sweep Extension
To use this extension for ITensor, one needs to use the packages ITensors, ITensorMPS and HDF5. As the set-up, a path needs to be created to store the result HDF5 files, and a method that converts OpSum and sites to MPO needs to be defined.
@@ 9 Practical Examples
FuzzifiED can help reproduce almost all the ED and DMRG results in the fuzzy-sphere works. We offer a series of such examples to help the users get started with the package. The code can be found in the examples directory of the source code repository.
@@
    defect_spectrum.jl calculates the spectrum of pinning-field defect in the 3D Ising model in $l_{z}=0,\mathscr{P}=\pm 1$ and $l_{z}=1$ sectors, calibrated by bulk $T$. This example reproduces Table I in Ref. [6].
@@
    defect_correlator.jl calculates the 1-pt function $\sigma$ and 2-pt function $\sigma\hat{\phi}$ of pinning-field defect in the 3D Ising model. The normalisation of the correlators requires extra bulk data. This example reproduces Figure 4 in Ref. [6].
@@
    defect_changing.jl calculates the spectrum of the defect-creation and changing operators of the pinning-field defect in the 3D Ising model. This example reproduces Table 2 and Figure 5 in Ref. [9].
@@
    defect_overlap.jl calculates the $g$-function of pinning-field defect in the 3D Ising model using the ovelaps between the bulk, defect ground state and the lowest defect-creation state. This example reproduces Figure 6 in Ref. [9].
@@
    cusp_dim.jl calculates the scaling dimension of the cusp of the pinning-field defect in the 3D Ising model as a function of the angle $\theta$. This example reproduces Table 2, upper panel in Ref. [11].
@@
    hsb_2l_spectrum.jl calculates the spectrum of $\mathrm{O}(3)$ Wilson-Fisher CFT using the bilayer Heisenberg model. This example reproduces Table I and Figure 2 in Ref. [8].
@@
    yang_lee_spectrum.jl calculates the spectrum of 3D non-unitary Yang-Lee CFT on the fuzzy sphere. This example reproduces Figure 11 in Ref. [22].
  25. 25.
    free_scalar_spectrum.jl calculates the spectrum of free real scalar. This example reproduces Figure 2 in Ref. [24].
  26. 26.
    free_scalar_correlator.jl calculates the correlator of $n^{i}$ for $\phi$, $\phi^{2}$ and $\pi$ of free real scalar. This example reproduces Figure 3 in Ref. [24].
  27. 27.
    o4_dqcp_spectrum.jl calculates the spectrum of $\mathrm{O}(4)$ DQCP on the fuzzy sphere. This example reproduces Figure 4 and Table I in Ref. [26].
  28. 28.
    su2_1_scal_spectrum.jl calculates the spectrum of the $\mathrm{U}(1)_{2}$ coupled to a complex scalar Chern-Simons matter CFT on the fuzzy sphere. This example reproduces Figures 2d and 3 in Ref. [27].
  29. 29.
    free_majorana_spectrum.jl calculates the spectrum of free Majorana fermion. This example reproduces Figure 5 in Ref. [29].
  30. 30.
    free_majorana_correlator.jl calculates the 2-pt correlator of free Majorana fermion. This example reproduces Figure 7 in Ref. [29].
  31. 31.
    bpf_220_spectrum.jl calculates the spectrum of the transition between bosonic Pfaffian and Halperin 220 described by gauged Majorana fermion CFT. This example reproduces Figure 4a in Ref. [30].
  32. 32.
    o3_wf_spectrum.jl calculates the spectrum of $\mathrm{O}(3)$ Wilson-Fisher CFT. This example takes the model from Ref. [32] and reproduces partly Tables I and II, Figures 1 and 2.
  33. 33.
    o2_wf_spectrum.jl calculates the spectrum of $\mathrm{O}(2)$ Wilson-Fisher CFT. This example reproduces Figure 3, upper panel in Ref. [33].
  34. 34.
    o2_free_spectrum.jl calculates the spectrum of $\mathrm{O}(2)$ free scalar CFT. This example reproduces Figure 4, upper-left panel in Ref. [33].
  35. 35.
    o4_wf_spectrum.jl calculates the spectrum of $\mathrm{O}(4)$ Wilson-Fisher CFT and demonstrates the general set-up for O(N) vector models. This example reproduces Figure 3, lower panel in Ref. [33].
  36. 36.
@@ Acknowledgements.
We extend our deepest gratitude to Yin-Chen He, who provided meticulous testing, unwavering support, and invaluable guidance throughout the development of the package. We also thank Andrew Liam Fitzpatrick, Wenhan Guo, Liangdong Hu, Xiang Li, Christian Voinea, Kay Jörg Wiese, Shuai Yang, and Wei Zhu for their suggestions and contributions to the package. Z. Z. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ A.1 Construction of Lin Table
In this appendix, we describe the construction of the Lin table used for a reverse look-up process that returns the index $i$ from the binary string $\mathit{conf}_{i}$ mentioned in Section 7.2.
Each binary string $\mathit{conf}_{i}$ is divided into a left part $l(\mathit{conf}_{i})$ from the $N_{or}$-th bit to the $(N_{o}-1)$-th bit and a right part $r(\mathit{conf}_{i})$ from the 0-th bit to the $(N_{or}-1)$-th bit from the right. The cut $N_{or}$ is typically taken as $N_{o}/2$. Two arrays $\mathit{lid}$ and $\mathit{rid}$ with length $2^{N_{o}-N_{or}}$ and $2^{N_{or}}$ are constructed, such that
$i=\mathit{lid}_{l(\mathit{conf}_{i})}+\mathit{rid}_{r(\mathit{conf}_{i})}.$ (83)
$c$ $\mathit{lid}_{c}$ $\mathit{rid}_{c}$ $i$ $\mathit{conf}_{i}$ $l_{i}$ $r_{i}$ $\mathit{lid}_{l_{i}}$ $\mathit{rid}_{r_{i}}$ $i$ $\mathit{conf}_{i}$ $l_{i}$ $r_{i}$ $\mathit{lid}_{l_{i}}$ $\mathit{rid}_{r_{i}}$
@@
Table 2: An example of explicit construction of the Lin table in the configurations on 6 sites filled with 3 particles. Here we use a shorthand notation $l_{i}=l(\mathit{conf}_{i})$ and $r_{i}=r(\mathit{conf}_{i})$
@@ A.2 Compressed Sparse Column (CSC) Sparse Matrix
In the CSC format, the matrix is stored in three arrays. We first index elements from $1$ to the number of elements $N_{\text{el}}$ in the ascending order from the column to column, and in each column from row to row.
@@
    The array $\mathit{colptr}$ records the column index of each element. It is of length $N+1$ where $N$ is the number of columns. The first $N$ elements record the index of the first element in each column ; the last element records $N_{\text{el}}+1$, so that the elements that belong to the $i$-th column have indices from $\mathit{colptr}_{i}$ to $\mathit{colptr}_{i+1}-1$.
@@
    The array $\mathit{rowid}$ records the row index of each element and is of length $N_{\text{el}}$.
@@
    The array $\mathit{elval}$ records the value of each element and is of length $N_{\text{el}}$.
@@
$\mathit{colptr}_{i}$ 1 4 7 8 11
$\mathit{rowid}_{i}$ 1 2 4 1 4 5 2 2 3 4
$\mathit{elval}_{i}$ 10 20 30 40 50 60 70 80 90 100
@@ A.3 Indexing the Boson Configurations
$\mathit{id}(n_{N_{o}}n_{N_{o}-1}\dots n_{1};N_{\max})$ $=N_{\textrm{cf}}(0)+N_{\textrm{cf}}(1)+\dots+N_{\textrm{cf}}(n_{N_{o}}-1)$
@@
$\mathit{id}(n_{N_{o}}n_{N_{o}-1}\dots n_{1};N_{\max})=\sum_{o=1}^{N_{o}}\sum_{i=0}^{n_{o}}\frac{\left[(o-1)+\left(N_{\max}-\sum_{o^{\prime}=o}^{N_{o}}n_{o^{\prime}}\right)\right]!}{(o-1)!\left(N_{\max}-\sum_{o^{\prime}=o}^{N_{o}}n_{o^{\prime}}\right)!}$ (87)
@@
$l(\{n_{o}\})=\mathit{id}(n_{N_{o}}n_{N_{o}-1}\dots n_{N_{or}+1};N_{\max})\quad\textrm{and}\quad r(\{n_{o}\})=\mathit{id}(n_{N_{or}}n_{N_{or}-1}\dots n_{1};N_{\max}).$ (88)
Appendix B Landau Levels and Pseudo-Potentials on the Torus
In this appendix, we describe the construction of Landau levels and pseudo-potentials on the torus [142, 143, 144], mentioned in Sections 4.4 and 7.11.
@@ B.1 Lowest Landau Level
The torus is nothing more than a $l_{x}\times l_{y}$ rectangle with periodic boundary conditions, i. e. $x\sim x+l_{x}$ and $y\sim y+l_{y}$. We adopt the Landau gauge $A_{x}=0,A_{y}=xB$ and set the magnetic length as the unit length $l_{B}=1$. In this way, the number of orbitals is $N_{m}=l_{x}l_{y}/2\pi$. The lowest Landau level wave-functions are labelled by the momenta along $y$ direction $m$
@@
$\psi_{f}^{\dagger}(\mathbf{r})=\phi_{m}(\mathbf{r})c^{\dagger}_{mf},\quad\mathbf{r}=(x,y)$ (90)
@@ B.2 Many-Body Interaction
$H_{\text{int}}=\frac{1}{N_{m}^{2}}\int\mathrm{d}^{2}\mathbf{r}_{1}\mathrm{d}^{2}\mathbf{r}_{2}\,U(\mathbf{r}_{1}-\mathbf{r}_{2})n(\mathbf{r}_{1})n(\mathbf{r}_{2}),\qquad n(\mathbf{r})=\psi^{\dagger}(\mathbf{r})\psi(\mathbf{r})$ (91)
@@
$H_{\text{int}}=\sum_{m_{1}m_{2}m_{3}m_{4}}\sum_{\mathbf{q}}\frac{U_{\mathbf{q}}}{2N_{m}}e^{iq_{x}(X_{m_{1}}-X_{m_{3}})}e^{-q^{2}/2}\delta^{\prime}_{t,m_{1}-m_{4}}\delta^{\prime}_{m_{1}+m_{2},m_{3}+m_{4}}c^{\dagger}_{m_{1}}c^{\dagger}_{m_{2}}c_{m_{3}}c_{m_{4}},$ (94)
@@
The interactions can alternatively be organised into pseudo-potentials $U_{l}$, describing the energy cost of a pair of electrons with relative angular momentum $l$.
@@
$H_{\text{int}}$ $=\sum_{lm_{1}m_{2}m_{3}m_{4}}U_{l}C_{m_{1}m_{2}m_{3}m_{4}}^{l}c^{\dagger}_{m_{1}}c^{\dagger}_{m_{2}}c_{m_{3}}c_{m_{4}}$
@@ 41 Es, sts = dmrg(hmt, sti_m ;
C.4 DMRG with Easy-Sweep